Two cars, A and B, travel in a straight line. The distance of A from the starting point is given

as a function of time by π‘₯𝐴(𝑑) = (2.60 π‘š / s) 𝑑 + (1.20π‘š /

𝑠2) 𝑑2. The distance of B from the starting point is

π‘₯𝐡(𝑑) = (2.80π‘š/𝑠2) 𝑑2 βˆ’ (0.20 π‘š/s3)t3 .

(a) Which car is ahead just after they leave the starting point?

(b) At what time(s) are the cars at the same point?

(c) At what time(s) is the distance from A

to B neither increasing nor decreasing?

(d) At what time(s) do A and B have the same

acceleration?


1
Expert's answer
2021-03-09T10:50:32-0500

a)


"v_{a0}=2.6\frac{m}{a}\\v_{b0}=0\frac{m}{a}"

Car A is ahead just after they leave the starting point

b)


"0.2t^2-1.6t+2.6=0\\t_1=0\ s\\t_2=2.27\ s\\t_3=5.73\ s"

c)


"0.6t^2-3.2t+2.6=0\\t_1=1\ s\\t_2=4.33\ s"

d)


"t=\frac{2.8-1.2}{3(0.2)}=2.67\ s"